Bill had 8600.00 in a savings account on April 1. He then deposited 800.00 on May 5 and an additional 350.00 on June 20. Find the balance in the account on July 1 if the accounts pays 4% compunded coninuosly.
I am really stuck on this one. Please help.

Answer by scott8148(6628)

you need to use the continuous compounding formula [A=Pe^(rt)] 3 times; and then sum the results

a) P=8600 __ t=91/365

b) P=800 __ t=57/365

c) P=350 __ t=11/365